OpenStudy (anonymous):

O(3^n), O(n^2),O(1), O(n log n) increasing order

OpenStudy (anonymous):

\(O(1)\) is evidently the smallest

ganeshie8 (ganeshie8):

O(3^n) is the biggest its obvious

OpenStudy (anonymous):

then \(\log(n)\) grows slower than \(n\) so \(O(n\log(n))\) is smaller than \(O(n^2)\)

hartnn (hartnn):

simpler method, put n as very large number, then arrange your answer in increasing order. that would be the order of O's

OpenStudy (anonymous):

thanks to all...... O(1)<O(n long n) < O(n^2)<O(3^n) Am I right???/

hartnn (hartnn):

yup :) good work.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

prove mathematical induction (sum 2^0 + 2^1 + ...+ 2^n is 2^n+1 - 1 for all n>=1

OpenStudy (anonymous):

how to prove???